OpenStudy (anonymous):

If the derivative of a function is y (the function it self), then f(x) = e^x. If the derivative of a function is y*y, what is f(x) ?

OpenStudy (anonymous):

so you want to integral of e^(2x)?

OpenStudy (anonymous):

\[\int{e^{2x}}dx=\frac{e^{2x}}{2}\]

OpenStudy (jamesj):

no, he means dy/dx = y^2

OpenStudy (anonymous):

so the function should be (x^3)/3 ?

OpenStudy (jamesj):

No, you see that doesn't work as y' = x^2.

OpenStudy (jamesj):

Hint: Note that for the equation dy/dx = y, we could solve using a method called separation of variables: \[\frac{dy}{y} = dx\] Integrating, ln y = x + C hence y = A e^x, where A = e^C Use the same method of separation of variables for dy/dx = y^2 @agd: don't write the answer; let him try and figure it out.

OpenStudy (anonymous):

Hmm I don't really follow, but dy/dx = y*y dy/y² = dx ln y² = x+ c y² = A e^x y = e^(c/2) + e^(x/2) ?

OpenStudy (jamesj):

No, you didn't integrate correctly: \[\int\limits \frac{dy}{y^2} = .... what?\]

OpenStudy (anonymous):

I don't follow, could you explain some more

OpenStudy (jamesj):

You said that integral is ln(y^2). That is incorrect. Let me ask you this. What is \[\frac{d \ }{dy} (1/y)\] ?

OpenStudy (anonymous):

d/dy 1/y = -1/y² d/dy 1/y2 -2/y³ ?

OpenStudy (jamesj):

Yes Hence the integral of 1/y^2 is ... what?

OpenStudy (jamesj):

\[\int\limits \frac{dy}{y^2}\]

OpenStudy (anonymous):

1/-y

OpenStudy (jamesj):

Yes. So plugging that back into the equation above with both x and y we have ... -1/y = x + c Hence y = ....

OpenStudy (anonymous):

y = -1/x+c

OpenStudy (jamesj):

Right. Let's set x = 0 for simplicity. Verify now directly be differentiating that \[\frac{dy}{dx} = y^2\]

OpenStudy (jamesj):

...by differentiating ...

OpenStudy (jamesj):

Sorry set C = 0

OpenStudy (jamesj):

sigh ... typos, sorry.

OpenStudy (jamesj):

So to clarify, the function we're now verifying does meet the equation dy/dx = y^2 is y(x) = -1/x

OpenStudy (anonymous):

You mean differentiating y = -1/x+c which comes to dy/dx -1 / x = 1/x² ... I'm doing something wrong here

OpenStudy (jamesj):

Cmon. If f(x) = x^n, what is df/dx?

OpenStudy (jamesj):

df/dx = n x^(n-1)

OpenStudy (anonymous):

n x^x-1

OpenStudy (anonymous):

n-1

OpenStudy (jamesj):

Hence here, with y = -1/x = - x^(-1), dy/dx = -(-1)x^(-1-1) = x^(-2) = 1/x^2 Now does that satisfy the equation dy/dx = y^2 ?

OpenStudy (anonymous):

nope

OpenStudy (jamesj):

YES, it does

OpenStudy (jamesj):

Because y^2 = (-1/x)^2 = 1/x^2

OpenStudy (jamesj):

Hence indeed dy/dx = y^2 when y = -1/x

OpenStudy (anonymous):

oh like that

OpenStudy (anonymous):

I read the whole equation wrong

OpenStudy (anonymous):

the equation says does the derivative of f(x) equate f(x)^2, right?

OpenStudy (jamesj):

Yes

OpenStudy (jamesj):

I.e., dy/dx = y^2

OpenStudy (jamesj):

Unless you tell me otherwise.

OpenStudy (anonymous):

Thanks man!

OpenStudy (jamesj):

Do yourself a favor: Take a blank piece of paper and write out the solution to the problem. When you can do that without looking at this website or the last time you wrote it out, then you know that you actually know it.

OpenStudy (anonymous):

yea thanks for the tip. Is that how you study ?

OpenStudy (jamesj):

Yes.